OpenStudy (anonymous):

graph the equation with a diameter that has endpoints at (-3,4) and (5,-2)

OpenStudy (goformit100):

@mary.rojas

OpenStudy (zzr0ck3r):

what would the center be?

OpenStudy (anonymous):

its doesnt say and i dont know how to find it

OpenStudy (zzr0ck3r):

its in the middle of the two points

OpenStudy (zzr0ck3r):

given points (a,b),(c,d) \[midpoint = (\frac{a+c}{2},\frac{b+d}{2})\]

OpenStudy (zzr0ck3r):

so whats the midpoint?

OpenStudy (anonymous):

1.1?

OpenStudy (zzr0ck3r):

im confused about the question

OpenStudy (zzr0ck3r):

is this the equation of a circle?

OpenStudy (zzr0ck3r):

how did you get 1.1?

OpenStudy (zzr0ck3r):

you shuold get a ordered pair

OpenStudy (anonymous):

??????I'm so long

OpenStudy (anonymous):

lost*

OpenStudy (anonymous):

It should be a circle. Centre = midpoint of the 2 given endpoints = ((5 - 3) / 2 , (4 - 2) / 2)) = (1 , 1) Radius = length of 1 endpoint to centre = sqrt ((5 - 1)^2 + (-2 - 1)^2) = 5 units (x - 1)^2 + (y - 1)^2 = 5^2 x^2 - 2x + 1 + y^2 - 2y + 1 = 25 x^2 + y^2 - 2x - 2y - 23 = 0

OpenStudy (zzr0ck3r):

you should get it to the standard form of a circle by "completing the square" \[x^2+y^2-2x-2y-23=0\\(x^2-2x)+(y^2-2y)-23=0\\((x-1)^2-1 )+((y-1)^2-1)-23=0\\(x-1)^2+(y-1)^2=5^2\]